# Measuring software innovation with open source software development data

### v1

@@ 4.1 Theoretical Implications
4.1.1 Making visible OSS innovation
@@
Our proposed measure of software innovation reduces a gap in the literature on innovation that has largely overlooked OSS outputs. Semantic versioning and the pronounced difference in downstream adoption of major versions compared to minor and patches may illustrate the concept of incremental versus radical innovation in a software context. This distinction is particularly important in the case of OSS, where contributions can range from small, incremental improvements to large, transformative updates.
@@
Additionally, in many emerging countries, measuring innovation is challenging due to limited data [74] or incentives to patent [20]. At the same time, OSS innovation holds significant potential in these regions due to its cost-effectiveness and accessibility, which can drive technological advancement without expensive proprietary solutions. Since traditional innovation metrics fall short in capturing the potential of OSS in these regions, our work represents an initial step toward developing new measurement methods, paving the way for further research that could enable country-level analyses and better assess OSS innovation, especially in areas where conventional metrics are inadequate.
@@
4.1.2 Complementing innovation indicators with granular, cost-effective, and real-time innovation measurement
@@
Established innovation indicators (discussed in § 2) face significant limitations in various contexts, particularly regarding timeliness and granularity. The Oslo Manual [2] highlights the challenge of measuring radical innovations due to delays in their effects, which can obscure the real-time dynamics of innovation. In contrast, our proposed measure enables more immediate evaluations of innovation–whether radical or incremental–even if allowing for a year to validate trends in dependents. Unlike patent applications, which are published only after an 18-month period from the earliest filing date, our methodology facilitates the analysis of how innovations are utilized, providing citation visibility akin to academic publications and practical implications similar to patents. Focusing on software innovation, our approach can also address insufficient sectoral and technological granularity of traditional metrics (e.g. with a focus on the ICT sector). Moreover, conventional indicators often incur high data collection costs, especially at scale. By leveraging accessible data from developer platforms such as GitHub, our methodology offers scalable access to data without substantial financial investment. Additionally, our approach employs established criteria for assessing software innovation, with only few qualitative aspects in comparison to e.g. survey-based innovation indicators. These qualitative aspects are primarily related to the documentation of release notes and the rigor of semantic versioning applied by contributors.
@@ 4.1.3 Transparency and visibility of individual contributions in the innovation ecosystem
Our measure of software innovation enables transparency in contributions: specific lines of code within an innovative package can be traced back to their creators. While not explored at length in this paper, this visibility could support further work to evaluate individual contributions to innovation over time and their relation to trends in academic publishing and other innovative fields. Building on this, social network analysis (SNA) presents an important avenue for future research in understanding the dynamics of OSS innovation ecosystems. Widely applied in innovation economics, SNA can offer a more in-depth exploration of how knowledge and innovations circulate within networks [75] As noted in the literature review, approaches to network analysis of OSS networks have already been undertaken (see, e.g. [56] or [57]). However, by incorporating dependency data and differentiating inventive steps through semantic versioning, SNA could provide a more granular map of interactions between contributors in OSS ecosystems and contribute to our understanding of how knowledge and innovation circulate within OSS projects. For example, this approach could reveal how collaborative efforts shape innovation, identifying key nodes or bottlenecks and informing policies aimed at fostering more effective knowledge exchange and innovation flow.
@@ 4.1.4 Potential self-fulfilling prophecy in release labeling and innovation adoption
Our measure may demonstrate a self-fulfilling prophecy that reflects the marketing of innovation. We find that the type of release is the strongest predictor of downstream adoption, and while it is (weakly) positively correlated with release complexity as assessed by release notes, developers may only pay attention to major releases due to their label or public sharing of a release in channels not observed in the GitHub data.This creates a risk of endogeneity: when a developer claims their new release is a major one, this assertion may spur downstream adoption based more on the label than the actual substance of the release.
This phenomenon parallels challenges seen in other innovation metrics, such as those for patents or R&D expenditures. In the patent landscape, the sheer number of filings does not inherently correlate with the quality or impact of the innovations. A high volume of patents might mask a lack of substantive progress, while truly innovative ideas could be underrepresented if they are not effectively communicated or marketed [76]. Similarly, with R&D expenditures, what is labeled as R&D can encompass a wide range of activities that do not necessarily lead to innovation; for instance, a significant portion of R&D spending might be directed towards marketing or administrative tasks rather than genuine innovative processes. This variability complicates the standardization of R&D as a measure of innovation [76].
In the context of our research, the label of a ”major release” could similarly drive adoption regardless of the technical merit or innovative substance of the release. This highlights the importance of considering how the communication surrounding releases can shape perceptions and behaviors in the OSS community. The influence of promotional language and activity may significantly affect how innovation is recognized and valued, leading to the potential misrepresentation of the innovation landscape. Further research on this possible phenomenon with off-platform data would support related research tracking promotional language and activity associated with academic research [77, 78].
@@ 4.1.5 Supporting comparisons of software language innovation ecosystems
Finally the measure of software innovation can be used to create a baseline by which adoption is expected for a major, minor, or patch version in a particular language. This could support research on language evolution and comparative analysis between programming language ecosystems, building on our observed variation between JavaScript, Python, and Ruby in growth of dependents two years after a release. Within a particular language, the expected baseline can also be used to identify especially innovative or impact releases early on, based on their divergent growth trajectory. Future work could describe this innovative subset and compare such a measure of ‘innovativeness’ to existing approaches like package citation combination novelty [55].
@@ 4.2.1 Informing national measurement and policy
Beyond the innovation literature, our measure of software innovation has practical implications for policymakers, practitioners, and researchers. The measure may be aggregated by jurisdiction, as has been done to-date with software development activities in the GitHub Innovation Graph and inform international comparisons of innovation [79]. Governments may look to the measure as one data point to inform the evaluation of national industrial strategies, technology-specific, e.g. AI, strategies, and other initiatives that aim to spur innovation. For practitioners in the private sector, our approach can support the evaluation of investment decisions and outcomes. By analyzing software innovation metrics, organizations can gain insights into market trends, competitive positioning, and potential return on investment. This can enable more informed decision-making regarding research and development initiatives, partnerships, and resource allocation.
4.2.2 Refining OSS innovation metrics
@@
For future research, we may consider the ratio of dependents to dependencies as a measure of a repository’s innovativeness, similar to the forward-to-backward citation ratio used in bibliometric and patent analysis [80]. This approach aligns with the concept of revealed preference, suggesting that when developers choose to depend on a project rather than writing code from scratch, they signal that the dependency represents a non-trivial inventive step. Our preliminary exploration of the top 1,000 GitHub repositories, ranked by stars, forks, and issue authors, has shown a positive correlation between the dependent-dependency ratio and the number of dependents. This suggests that repositories with a high ratio of dependents to dependencies can be considered more innovative, as they exert a greater influence on the OSS ecosystem relative to their reliance on other projects. Using major version releases as output units of innovation provides a clearer view of substantial changes in software development that could potentially have wider economic and technological impacts. While directly linking these releases to specific economic outcomes—such as increased demand for technical services, job creation, or shifts in industry productivity—remains subject to further research, this measure may serve as an early indicator of such impacts. For example, downstream adoption patterns following a major release could signal broader trends in the uptake of new technologies or the influence of OSS on industry innovation.
However, in this work, we chose not to rely on dependency data as the primary unit of innovation, as it appeared less useful for comparing individual releases within a project. Outside of major releases, projects rarely change their dependencies significantly, making this ratio less informative for release-to-release comparisons. Nonetheless, project-to-project comparisons using the dependent-dependency ratio remain a promising avenue for identifying innovative software. Future studies should further explore these methods to connect the proposed OSS innovation metrics to broader economic or sectoral trends.
While our measure of OSS innovation is primarily designed to track innovation within the software development ecosystem itself, future research could investigate potential links to broader outcomes, such as industry adoption rates or collaborative dynamics across programming communities. In this sense, the measure provides a foundation for more granular analysis of OSS innovation, without yet venturing into economic impact analysis.
@@ 4.3 Limitations
This work faces the following limitations. First, we analyze only direct dependents of a given package. However, this undervalues packages that have become central to software ecosystems, where their direct dependencies are in turn relied upon widely. Future work could seek to remedy this limitation by analyzing indirect dependencies, albeit at significantly higher computational costs.
Second, our analysis does not account for the role of software licenses in shaping innovation patterns. Different license types (e.g. GPL vs. Apache v2) may systematically influence both developer choices and dependency relationships, as licenses can either enable or constrain the ways packages build upon each other. Furthermore, we do not address the geographic dimension of software innovation diffusion, which could reveal important regional variations in OSS adoption and development patterns.
Third, mature software packages often have their own platforms for communicating and distributing releases that are not captured in GitHub data. For example, some version releases on GitHub may use release notes that link to further detail off-platform. Thus, our method may not accurately account for the complexity of mature packages.
Fourth, and relatedly, our analysis of release complexity examines the notes describing the release only. Further analysis could examine the code associated with each release directly as another, perhaps more direct, indicator of complexity. Fourth, we examine only three software languages. Future work could expand the language coverage of the measure. However, some languages, such as C and C++, do not have centralized package ecosystems or rely on GitHub’s Releases feature, which undermines the universal application of our measure.
Finally, while our use of LLMs to assess semantic content offers novel analytical possibilities, this approach has inherent limitations. Recent research suggests that LLM responses can vary significantly even with identical prompts [81, 82]. Future work could enhance reliability by implementing multiple prompt iterations and employing modal or mean responses as final classifications, alongside expanded human validation procedures.
@@ 5 Conclusion
This paper has introduced a novel method for measuring innovation in software, complementing traditional indicators such as scientific publications, patents, and standards. By analyzing GitHub data on package releases and dependency adoption, we make three empirical contributions. First, we observed that major releases are associated with significantly larger increases in package dependencies compared to minor or patch releases across JavaScript, Python, and Ruby ecosystems. Second, this trend holds true regardless of package ecosystem, initial package utilization, or time frame examined, up to two years post-release. Third, we found that release complexity, as rated by a prompted GPT-4 model, correlates slightly positively with dependency growth, particularly for JavaScript and Python packages. These findings offer new insights into OSS innovation dynamics and raise important questions for future research and discussion.

### v2

@@ 5.1 Theoretical Implications
5.1.1 Making Visible OSS Innovation and Its Nuances
@@
Our proposed measure of software innovation helps closing a critical gap in the innovation literature, which has largely overlooked open-source software (OSS) outputs. The distinct downstream adoption patterns tied to semantic versioning—major, minor, and patch releases—offer a compelling illustration of incremental versus radical innovation within a software context. This distinction is especially critical for OSS, whose contributions span small, iterative improvements to significant, transformative updates.
Crucially, our findings demonstrate that the impact of a release type on dependency growth is not uniform. Instead, it is highly contingent on the specific ecosystem (e.g., JavaScript, Python, Ruby), the package’s pre-existing popularity, and its version series maturity (e.g., zero-version, one-version, two-plus-version). For instance, while minor releases may show consistent benefits across some ecosystems and package sizes, major or patch releases might exhibit varying impacts depending on ecosystem culture or baseline popularity. This highlights that ”major” does not universally equate to ”most impactful” in terms of rapid adoption.
@@
Furthermore, in many emerging countries, measuring innovation is challenging due to limited data [74] or a lack of incentives to patent innovations [20]. OSS innovation, with its cost-effectiveness and accessibility, offers substantial potential in these regions to drive technological advancement without relying on expensive proprietary solutions. Since traditional innovation metrics often fail to capture OSS’s impact in these areas, our work provides an initial step toward developing new measurement methods. This paves the way for future research that could enable country-level analyses and more accurately assess OSS innovation, particularly where conventional metrics are inadequate.
@@
5.1.2 Complementing Innovation Indicators with Granular, Cost-Effective, and Context-Aware Measurement
@@
Established innovation indicators (discussed in § 2) face significant limitations, particularly regarding timeliness and granularity. The Oslo Manual [2] highlights the challenge of measuring radical innovations due to delays in their effects, which can obscure the real-time dynamics of innovation. In contrast, our proposed measure enables more rapid evaluations of innovation—radical or incremental—even with a one-year window to validate trends in dependents. Unlike patent applications, which are published only after an 18-month period from the earliest filing date, our methodology facilitates analyzing how innovations are utilized, providing citation visibility akin to scientific publications and practical implications similar to patents. By focusing on software innovation, our approach can also address the insufficient sectoral and technological granularity of traditional metrics (e.g., with a focus on the ICT sector). Moreover, conventional indicators often incur high data collection costs, especially at scale. By leveraging accessible data from developer platforms such as GitHub, our methodology offers scalable access to data without substantial financial investment. Additionally, our approach employs established criteria for assessing software innovation, with few qualitative aspects compared to, for example, survey-based innovation indicators. These qualitative aspects primarily relate to the documentation of release notes and the rigor of semantic versioning applied by contributors.
@@ 5.1.3 Transparency and Visibility of Individual Contributions in the Innovation Ecosystem
Our measure of software innovation enhances transparency in contributions: specific lines of code within an innovative package can be traced back to their creators. While not extensively explored in this paper, this visibility could support further work to evaluate individual contributions to innovation over time and their relation to trends in academic publishing and other innovative fields. Building on this, social network analysis (SNA) presents an important avenue for future research in understanding the dynamics of OSS innovation ecosystems. Widely applied in innovation economics, SNA can offer a more in-depth exploration of how knowledge and innovations circulate within networks [75]. As noted in the literature review, approaches to network analysis of OSS networks have already been undertaken (see, e.g., [56] or [57]). However, by incorporating dependency data and differentiating inventive steps through semantic versioning, SNA could provide a more granular map of interactions between contributors in OSS ecosystems and contribute to our understanding of how knowledge and innovation circulate within OSS projects. For example, this approach could reveal how collaborative efforts shape innovation, identifying key nodes or bottlenecks and informing policies aimed at fostering more effective knowledge exchange and innovation flow.
@@ 5.1.4 Potential Self-Fulfilling Prophecy in Release Labeling and Innovation Adoption
Our measurements may demonstrate a self-fulfilling prophecy that reflects the marketing of innovation. We find that the type of release is a strong predictor of downstream adoption. While there is a weak positive correlation between semantic versioning and release complexity, as assessed by release notes, developers may primarily pay attention to major releases due to their label or public sharing of a release in channels not observed in the GitHub data. This creates a risk of endogeneity: when a developer claims their new release is a major one, this assertion may spur downstream adoption based more on the label than the actual substance of the release.
This phenomenon parallels challenges in other innovation metrics, such as those for patents or R&D expenditures. In the patent landscape, the sheer number of filings does not inherently correlate with the quality or impact of the innovations. A high volume of patents might mask a lack of substantive progress, while truly innovative ideas could be underrepresented if they are not effectively communicated or marketed [76]. Similarly, with R&D expenditures, what is labeled as R&D can encompass a wide range of activities that do not necessarily lead to innovation; for instance, a significant portion of R&D spending might be directed toward marketing or administrative tasks rather than genuine innovative processes. This variability complicates the standardization of R&D as a measure of innovation [76].
In the context of our research, the label of a ”major release” could similarly drive adoption regardless of the release’s technical merit or innovative substance. This highlights the importance of considering how the communication surrounding releases can shape perceptions and behaviors in the OSS community. The influence of promotional language and activity may significantly affect how innovation is recognized and valued, potentially leading to a misrepresentation of the innovation landscape. Further research on this possible phenomenon with off-platform data would support related research tracking promotional language and activity associated with academic research [77, 78].
@@ 5.1.5 Supporting Comparisons of Software Language Innovation Ecosystems
Finally, our measurements of software innovation can be used to create a baseline for expected adoption for major, minor, or patch versions within a particular ecosystem, package size, and version series set. This could support research on language evolution and comparative analysis between programming language ecosystems, building on our observed variation between JavaScript, Python, and Ruby in dependent growth one year after a release. Within a particular language, the expected baseline can also identify especially innovative or impactful releases early on based on their divergent growth trajectory. Future work could describe this innovative subset and compare such a measure of ’innovativeness’ to existing approaches like package citation combination novelty [55].
@@ 5.2.1 Informing National Measurement and Policy
Beyond the innovation literature, our measure of software innovation has practical implications for policymakers, practitioners, and researchers. Measurements may be aggregated by jurisdiction, as has been done to date with software development activities in the GitHub Innovation Graph, and inform international comparisons of innovation [79]. Governments may look to the measure as one data point to inform the evaluation of national industrial strategies, technology-specific (e.g., AI) strategies, and other initiatives that aim to spur innovation.
For practitioners in the private sector, our approach can support the evaluation of investment decisions and outcomes. By analyzing software innovation metrics, organizations can gain insights into market trends, competitive positioning, and potential return on investment. This can enable more informed decision-making regarding research and development initiatives, partnerships, and resource allocation.
5.2.2 Refining OSS Innovation Metrics with Contextual Normalization
@@
Our findings strongly suggest that future OSS innovation and adoption measures should incorporate contextual normalization schemes*, akin to ”Field Weighted Citation Impact” used in scientometrics [80]. As research impact is normalized by field and publication year, OSS innovation metrics should account for the specific ecosystem, package size, and version series. This ensures that the measured ”impact” of a release (e.g., in terms of dependent growth) is evaluated relative to what is expected within its specific context, providing a more accurate and equitable assessment of its innovativeness.
For future research, we may consider the ratio of dependents to dependencies as a measure of a repository’s innovativeness, similar to the forward-to-backward citation ratio used in bibliometric and patent analysis [81]. This approach aligns with the concept of revealed preference, suggesting that when developers choose to depend on a project rather than writing code from scratch, they signal that the dependency represents a non-trivial inventive step. Our preliminary exploration of the top 1,000 GitHub repositories, ranked by stars, forks, and issue authors, has shown a positive correlation between the dependent-dependency ratio and the number of dependents. This suggests that repositories with a high ratio of dependents to dependencies can be considered more innovative, as they exert a greater influence on the OSS ecosystem relative to their reliance on other projects.
Using semantic versioning as output units of innovation provides a clearer view of substantial changes in software development that could potentially have broader economic and technological impacts. While directly linking these releases to specific economic outcomes—such as increased demand for technical services, job creation, or shifts in industry productivity—remains subject to further research, this measure may serve as an early indicator of such impacts. For example, downstream adoption patterns following a major release could signal broader trends in the uptake of new technologies or the influence of OSS on industry innovation.
However, in this work, we chose not to rely on a project’s ratio of dependents to dependencies as the primary unit for *individual release* innovation, as it appeared less helpful in comparing individual releases within a project due to infrequent significant dependency changes. Nonetheless, project-to-project comparisons using the dependent-dependency ratio remain promising for identifying innovative software. Future studies should explore these methods to connect the proposed OSS innovation metrics to broader economic or sectoral trends.
While our measure of OSS innovation is primarily designed to track innovation within the software development ecosystem itself, future research could investigate potential links to broader outcomes, such as industry adoption rates or collaborative dynamics across programming communities. In this sense, our log-difference measurements provide a foundation for more granular OSS innovation analysis without venturing into economic impact analysis.
@@ 5.3 Limitations
This work faces the following limitations. First, we analyze only the direct dependents of a given package. This undervalues packages that have become central to software ecosystems, where their direct dependencies are, in turn, relied upon widely. Future work could remedy this limitation by analyzing indirect dependencies, albeit at significantly higher computational costs.
Second, our analysis does not account for the role of software licenses in shaping innovation patterns. Different license types (e.g., GPL vs. Apache v2) may systematically influence developer choices and dependency relationships, as licenses can enable or constrain how packages build upon each other. Furthermore, we do not address the geographic dimension of software innovation diffusion, which could reveal important regional variations in OSS adoption and development patterns.
Third, mature software packages often have platforms for communicating and distributing releases that are not captured in GitHub data. For example, some version releases on GitHub may use release notes that link to further details off-platform. Thus, our method may not accurately account for the complexity of mature packages.
Fourth and relatedly, our analysis of release complexity examines only the notes describing the release. Further study could examine the code associated with each release directly as another, perhaps more direct, indicator of complexity. Fifth, we examine only three software languages. Future work could expand the language coverage of the measure. However, some languages, such as C and C++, do not have centralized package ecosystems or rely heavily on GitHub’s Release feature, which undermines the universal application of our measure.
Finally, while using LLMs to assess semantic content offers novel analytical possibilities, this approach has inherent limitations. Recent research suggests that LLM responses can vary significantly even with identical prompts [82, 83]. Future work could enhance reliability by implementing multiple prompt iterations and employing mean responses as final classifications alongside expanded human validation procedures.
@@ 6 Conclusion
This paper introduced a method for measuring software innovation, complementing traditional indicators by analyzing GitHub data on package releases and dependency adoption. Our findings reveal that the impact of different release types (major, minor, patch) on dependency growth depends on factors like the programming language ecosystem, package popularity, and version series maturity. This means a ”major” release isn’t always the most impactful; its effect depends on its context. Interestingly, while maintainers generally align a release’s technical complexity with its semantic versioning type, the complexity of a release shows limited direct correlation with its downstream dependency growth. This suggests that release type can act as a signaling mechanism, influencing adoption somewhat independently of the technical effort involved. For future research, this highlights the need for normalized approaches to measuring OSS innovation, similar to ”Field Weighted Citation Impact,” to account for these contextual nuances and provide a more accurate assessment of a release’s impact.
@@ Appendix A Complexity rating LLM prompt
Appendix B One-Year 90 Day Timepoint Distributions of Log-Differences of Dependents for Pre-Existing Package Utilization Bins
Figure 5: Distributions of log-difference dependents split by Ecosystem (NPM, PyPI, and RubyGems), pre-release package size (small, medium, large, and huge), release type (major, minor, or patch), and over multiple 90-day timepoints for one year after release.
Appendix C Two-Year 180 Day Timepoint Distributions of Log-Differences of Dependents for Pre-Existing Package Utilization Bins
Figure 6: Distributions of log-difference dependents split by Ecosystem (NPM, PyPI, and RubyGems), pre-release package size (small, medium, large, and huge), release type (major, minor, or patch), and over multiple 180-day timepoints for two years after release.
Appendix D One-Year 90 Day Timepoint Distributions of Log-Differences of Dependents for Version Series Bins
Figure 7: Distributions of log-difference dependents split by Ecosystem (NPM, PyPI, and RubyGems), version series (zero-version, one-version, and two-plus-version), release type (major, minor, or patch), and over multiple 90-day timepoints for one year after release.
Appendix E Two-Year 180 Day Timepoint Distributions of Log-Differences of Dependents for Version Series Bins
Figure 8: Distributions of log-difference dependents split by Ecosystem (NPM, PyPI, and RubyGems), version series (zero-version, one-version, and two-plus-version), release type (major, minor, or patch), and over multiple 180-day timepoints for two years after release.